OpenStudy (annie96):

Determine the slope of the line that contains the points (0, 2) and (-5, 1). (Points : 1) 1/5 5/2 -1/5 5

OpenStudy (luigi0210):

Use the point slope formula

OpenStudy (anonymous):

the answer is 1/-5

OpenStudy (anonymous):

\[\frac{ 1-2 }{-5-0 }=\frac{ 1 }{5 }\]